Two trains made the same 350-mile run. Since one train traveled 15 mph faster than the other, it arrived 3 hours earlier. Find the speed of each train.
This is how I set it up, but I'm lost after that
350/x +3= 350/x-15

Answer by umnikmo(75)

V1=V2+15
t1=350/V1=350/(V2+15)
t2=350/V2
t1+3=t2
350/(V2+15)+3=350/V2
multiplying everything by (V2+15)V2
350V2+3V2^2+45V2=350V2+5250
3V2^2+45V2^2-5250=0
Solution is : V2=35mph
V1=35+15=50mph